OpenStudy (anonymous):

cos(x)-sin(x)=sqrt(2)sin(x) prove that: cos(x)+sin(x)=sqrt(2)cos(x)

OpenStudy (anonymous):

1.f(x)= log(\[\sqrt{x ^{2}+1}\]-x) 2.f(x)= xlog( x+\[\sqrt{x ^{2}+1}\] ) find which is odd and which is even